OpenStudy (anonymous):

Find the area of a rectangle with length of 6 inches and width of 2 feet. 12 in.² 72 in.² 144 in.²

OpenStudy (tennistar):

12 inches 6*2=12 Area is the length times the width

jimthompson5910 (jim_thompson5910):

It's not 6*2=12 because you need to convert everything to the same units first.

jimthompson5910 (jim_thompson5910):

Keep in mind that 1 ft = 12 inches, so 2 ft = 2*12 = 24 inches

OpenStudy (tennistar):

oh wait, yeah I didnt see inches and feet

OpenStudy (tennistar):

so 144, sorry

OpenStudy (anonymous):

ohhh, i get it.(: thank you!

jimthompson5910 (jim_thompson5910):

so it's best to always watch out for that because they threw in that trap that a lot of people fall for

OpenStudy (tennistar):

ok